# Neutron Tagging Can Greatly Reduce Spallation Backgrounds in Super-Kamiokande

### v1

@@ I Introduction
The ability to use neutron tagging to identify potential isotope-producing showers depends on the neutron detection efficiency. We focus on the present Super-K with gadolinium, showing that it can greatly reduce spallation backgrounds relative to the previous Super-K with pure water. This highlights the importance of considering added gadolinium for Hyper-Kamiokande (Hyper-K) [34, 35], which will be at a shallower depth, where the flux of muons is about five times higher [35]. Further, the physical insight that isotope production often happens in hadronic showers with multiple neutrons will also be important for scintillator detectors like JUNO [36, 37, 38], which is also at a shallower depth. As detailed in the Appendices, our insight significantly generalizes upon a single-neutron technique used to reject 11C backgrounds in scintillator detectors [39, 40, 41, 42].
@@ II.1 Exposure to cosmic-ray muons
Despite the rock overburden [11, 12], Super-K is exposed to cosmic-ray muons at a rate of $\sim$2 s-1 [24]. At the depth of Super-K, the muons have an average energy of 271 GeV (almost all throughgoing), with most of the distribution being in the energy range 30–700 GeV [43]. Isotope production rates are only moderately sensitive to the detailed shape of Super-K’s muon spectrum. Most muons are relatively downgoing, with their angular distribution peaking at vertical and with $\sim$80of them being within $60^{\circ}$ of vertical. Taking into account the variations in the muon entry points and angles, the average muon track length inside the fiducial volume is $\sim$24 m [44] (for comparison, the longest diagonal path is $\sim$44 m). The majority ($\sim$89%) of muons arrive as individual throughgoing particles with long pathlengths, but $\sim$7of them arrive in bundles with parallel tracks due to being generated by the same primary cosmic ray, and $\sim$4are corner-clipping muons with small track lengths, usually less than $\sim$7 m.
@@ II.2 Muon propagation in water
where $X$ is the mass column density ($\rho L$, the product of the mass density and the path length, which is quoted in units of g cm-2), and where $\alpha(E)$ and $\beta(E)$ are the ionization and radiation coefficients, respectively. The energy dependence of $\alpha(E)$ and $\beta(E)$ is modest; the majority of the energy dependence in Eq. (1) arises from the explicit $E$ factor in the second term. The muon critical energy, $E_{c}$, is the energy at which the ionization and radiation losses are equal. In water, this is 1.03 TeV [54]. At the average energy of 271 GeV, ionization losses are dominant but radiation losses are appreciable (and cause the most relevant showers).
@@
The $\beta(E)$ term accounts for radiation energy losses through interactions with atomic nuclei. The most relevant processes here are pair production, bremsstrahlung, and photonuclear interactions. Pair-production losses are almost continuous (with small fluctuations), with an average rate of $\sim$0.4 MeV g-1 cm2, so they contribute to the minimum energy loss of muons crossing the detector. The other processes have much larger fluctuations, often yielding energetic daughter particles that can, in turn, initiate electromagnetic or hadronic showers. Importantly, the dominant energy losses of muons are due to electromagnetic processes, so electromagnetic showers are common and hadronic showers are rare.
@@ II.3 Showers induced by muons
The length scale of particle showers is determined by the mean free path of their underlying processes in the medium they transverse [59, 60, 55]. For electromagnetic showers in water, the mean free path of bremsstrahlung and pair production processes, or the radiation length $X_{0}$, is about 36 cm [55]. The distance to the shower’s maximum is $X_{0}\ln(E_{0}/E_{c})$, where $E_{0}$ is the initial energy of the first particle starting the shower. The full length of the shower is several times larger [61]. For hadronic showers, the mean free path for pion interactions is numerically somewhat larger than $X_{0}$, but the multiplicity of the particles in each generation is larger than in electromagnetic showers, leading to a comparable length for the showers. In Super-K, muon-induced electromagnetic and hadronic showers are less than 10 meters long even at the highest relevant energies, so they are often fully contained within the fiducial volume (when they are not, they can be treated as showers of lower energy).
Electromagnetic showers are quite narrow, with a characteristic (Molière) radius of $\sim$0.1 m in water [55], while hadronic showers, which have larger fluctuations at a given energy, have a characteristic radius of $\sim$1 m, about twice the hadronic interaction length. While many of the secondary particles in both types of showers are collimated in the direction of the muon, lower-energy secondaries have a broader range of directions, which is important for shower reconstruction [32].
@@ II.4 Production of isotopes by showers
The total path length of the secondary particles in a shower is directly related to its initial energy because the number of secondary particles in a shower grows linearly with shower energy. As noted, the total energy in showers can be estimated from the total amount of Cherenkov light produced by the muon, minus the expectation from the non-showering energy loss components for the muon. Therefore, the larger the light yield of a shower, the greater the cumulative path length of its secondary particles, which leads to a higher probability of producing an isotope.
@@ II.5 Super-K cuts on spallation isotopes
This improved likelihood cut was used in the latest solar neutrino analysis by Super-K [23]. In addition to that, they used a preselection cut that targets clusters of two or more isotopes produced by the same muon, which they called a multiple spallation cut. In principle, this is equivalent to imposing cuts on muons with large showers, as advocated for in Ref. [31], where it was noted that these could lead to multiple isotopes. However, their multiple spallation cut has the technical advantage of not requiring muon reconstruction. Overall, the latest Super-K analysis achieved an efficiency of $\sim$90with a 10.5% deadtime [25], a significant improvement over earlier work by Super-K. For the data period with neutron tagging in pure water, they also applied cuts based on clouds of neutrons (two or more detected, which implies $\sim$10 times more produced) [50, 51, 23]. These cuts used geometric estimates of the shower size and position based on the multiplicity and locations of the neutron captures. The impact of these geometric cuts was limited by the low neutron statistics. We advise that it is better to estimate the shower size and position based on its abundant Cherenkov light. This neutron cloud technique further reduced the deadtime to 8.6% [23, 25].
@@ III.1 Average neutron yields
Neutrons, which are abundantly produced by muons in underground experiments, can cause a wide variety of difficult-to-reject detector backgrounds, not only through spallation decays. This is why neutrons are sometimes called “the rats of the lab.” There have been many measurements of neutron yields, but they have typically been reported as a function of the average energy of the muons at a given depth; see, e.g., Refs. [72, 49, 73, 44]. This is not a very useful physical description. Instead, it would be better to measure the neutron yield as a function of shower energy, as this would put measurements at different depths on a common basis closely related to the underlying physics. As noted above, the shower energy can be determined from the measured total energy loss in the detector minus the predicted minimum energy loss. (At energies below the critical energy, it would be more accurate to say injected energy, as proper showers will not develop.) The neutron yield measurements would be even more informative if they were measured separately for hadronic and electromagnetic showers.
@@
In Appendix A, we give further details on the correlated yields of final-state neutrons, pions (and hence muons), and spallation isotopes, showing that they are all significantly higher for hadronic shower than electromagnetic showers (Fig. A.1).
@@ IV.2 Neutron yields of hadronic showers are large
Figure 5 shows that hadronic showers almost always produce neutrons. As in Sec. III, we define $N_{n}$ as the number of low-energy neutrons left at the end of a shower; this is the same number that will undergo capture, for which the detection efficiency depends on the gadolinium concentration. The number of neutrons produced is often large (as shown above in Fig. 3). If we remake Fig. 5, instead requiring, for example, $N_{n}\geq 3$, the results only change slightly. The fraction of hadronic showers that do not produce any neutrons is only 3%, and the majority of these are low-energy hadronic showers or electromagnetic showers with small hadronic components, which are less efficient at producing isotopes anyway. These points mean that neutrons are extremely efficient at identifying hadronic showers, which is good.
[…]
IV.3 Isotopes are produced with neutrons
@@ IV.4 Neutron-producing muons are rare
Figure 11 — another of our main results — shows the fraction of muons with neutron yields at or above a given value, assuming different tagging efficiencies. These values represent the fraction of the integral contributed by the bins in Fig. 7 above different $N_{n}$ thresholds. The sharp drop in the fraction of muons from $N_{n}=0$ to $N_{n}=1$ shows that about 89% of muons that cross Super-K produce no neutrons at all. This is because $\sim$97of muon-induced showers are electromagnetic (Fig. 4) and only $\sim$9of those electromagnetic showers produce neutrons (Fig. 6). The rate of all muons is $\sim$2 s-1, so the rate of neutron-producing muons is only $\sim$0.2 s-1.
@@ IV.5 Recommendations for implementation
Figure 12: Deadtime and efficiency comparison for cylinder cuts based on shower energy (blue solid line) and neutron yields (dotted green line). The blue and green points represent different thresholds for $E_{\text{sh}}$ and $N_{n}$ with increments of 1 GeV and 1 neutron, respectively. The grey band highlights the efficiency improvement at $\sim$10deadtime.
@@
Figure 12 (before case) shows the deadtime and background-rejection efficiency for cylinder cuts where we vary the threshold for $E_{\text{sh}}$, meaning that showers above a certain energy are always rejected. These cuts can approximate (for a cut-off between 6–7 GeV) the performance of the current likelihood used by Super-K Super-K, which rejects $\sim$90of the backgrounds with $\sim$10deadtime.
@@ Appendix A Final-state yields from showers
Figure A.1: Numbers of neutron captures, muon decays, and isotopes left after hadronic and electromagnetic showers as a function of their energy. The differences between the hadronic and electromagnetic showers are large.
@@
Figure A.1 shows the neutron, muon, and isotope yields in hadronic and electromagnetic showers as a function of the shower energy. The main feature is that the yields are approximately one order of magnitude lower in the electromagnetic case. In the hadronic case at the lowest energies, the results are significantly affected by the fact that half of the pions we inject are negative. In scintillator detectors, the muon and neutron yields are comparable, but the isotope yield is much higher, principally due to 11C (see Apeendix. C).
@@
Figure A.2: Neutron yield distribution for muons at Super-K, similar to Fig. 7, but extending to larger neutron yields and using bin widths of of $N_{n}=100$ instead of 1.
@@ Appendix B Spatial profiles
Figure B.1: Average longitudinal profiles of neutron captures, isotope production locations, and electron stopping points in 10-GeV hadronic and electromagnetic showers. The differences between hadronic and electromagnetic showers are small.
@@
Figure B.1 shows the average longitudinal profiles of neutron captures, isotope production, and electron stopping points in 10-GeV electromagnetic and hadronic showers. Both have a longitudinal extent of $\sim$6 m, and the shapes of their longitudinal profiles have small differences. One of the main differences is that the neutron and isotope profiles in hadronic showers peak at a smaller distance. This is because the initial interactions of the injected pions can produce neutrons and isotopes.
Figure B.2: Average lateral profiles of neutron captures, relevant isotope production locations, and electron stopping points in 10-GeV hadronic and electromagnetic showers. The differences between hadronic and electromagnetic showers are large.
Figure B.2 shows the average lateral profiles of neutron captures, isotope production, and electron stopping points around the tracks of 10-GeV electromagnetic and hadronic showers. The area factor $2\pi rdr$ is included. Compared to the longitudinal profiles, the lateral profiles of electromagnetic showers and hadronic showers have large differences.
For electromagnetic showers, the lateral extent is characterized by the Molière radius ($\sim$10 cm in water). The majority of the shower energy is deposited within 0.5 m around the shower track, as seen in the bottom panel of Fig. B.2. At 10 GeV, electromagnetic showers have an average neutron yield of $\sim$0.6. Neutrons at energies above tens of MeV can travel far away from the track because they undergo many scattering interactions before capturing, unlike thermal neutrons, but most of them still capture within $\sim$1.5 m from the shower. Isotope production in electromagnetic showers is mainly caused by these neutrons, so the isotopes are also contained within $\sim$1.5 m, but there is a larger peak at close distances to the shower because of isotopes produced by gamma rays directly or by neutrons that interact quickly.

### v2

@@ I Introduction
The ability to use neutron tagging to identify potential isotope-producing showers depends on the neutron detection efficiency. We focus on the present Super-K with gadolinium, showing that it can greatly reduce spallation backgrounds relative to the previous Super-K with pure water. This highlights the importance of considering added gadolinium for Hyper-Kamiokande (Hyper-K) [34, 35], which will be at a shallower depth, where the flux of muons is about five times higher [35]. Further, the physical insight that isotope production often happens in hadronic showers with multiple neutrons will also be important for scintillator detectors like JUNO [36, 37, 38], which is also at a shallower depth. As detailed in the Appendices, our insights significantly improve upon a standard technique (TFC, or three-fold coincidence between muon, neutron, and candidate isotope) used to reject 11C in scintillator detectors  [39, 40]. It also generalizes upon new approaches used by Borexino and KamLAND-Zen that incorporate neutron multiplicity to tag 11C and xenon spallation events [41, 42].
@@ II.1 Exposure to cosmic-ray muons
Despite the rock overburden [11, 12], Super-K is exposed to cosmic-ray muons at a rate of $\sim$2 s-1 [24]. At the depth of Super-K, the muons have an average energy of 271 GeV (almost all throughgoing), with most of the distribution being in the energy range 30–700 GeV [43]. Isotope production rates are only moderately sensitive to the detailed shape of Super-K’s muon spectrum. Most muons are relatively downgoing, with their angular distribution peaking at vertical and with $\sim$80of them being within $60^{\circ}$ of vertical. Taking into account the variations in the muon entry points and angles, the average muon track length inside the fiducial volume is $\sim$24 m [44] (for comparison, the longest diagonal path is $\sim$44 m). The majority ($\sim$89%) of muons arrive as individual throughgoing particles with long pathlengths, but $\sim$7of them arrive in bundles with parallel tracks due to being generated by the same primary cosmic ray (described further in Sec. II.5), and $\sim$4are corner-clipping muons with small track lengths, usually less than $\sim$7 m.
@@ II.2 Muon propagation in water
where $X$ is the mass column density ($\rho L$, the product of the mass density and the path length, which is quoted in units of g cm-2), and where $\alpha(E)$ and $\beta(E)$ are the ionization and radiation coefficients, respectively. The energy dependence of $\alpha(E)$ and $\beta(E)$ is modest; the majority of the energy dependence in Eq. (1) arises from the explicit $E$ factor in the second term. The muon critical energy, $E_{c}$, is the energy at which the ionization and radiation losses are equal. In water, this is 1.03 TeV [54, 56]. At the average energy of 271 GeV, ionization losses are dominant but radiation losses are appreciable (and cause the most relevant showers).
@@
The $\beta(E)$ term accounts for radiation energy losses through interactions with atomic nuclei. The most relevant processes here are pair production, bremsstrahlung, and photonuclear interactions. Pair-production losses are almost continuous (with small fluctuations), with an average rate of $\sim$0.4 MeV g-1 cm2, so they contribute to the minimum energy loss of muons crossing the detector. The other processes have much larger fluctuations, often yielding energetic daughter particles that can, in turn, initiate electromagnetic or hadronic showers [56]. Importantly, the dominant energy losses of muons are due to electromagnetic processes, so electromagnetic showers are common and hadronic showers are rare.
@@ II.3 Showers induced by muons
The length scale of particle showers is determined by the mean free path of their underlying processes in the medium they transverse [60, 61, 55]. For electromagnetic showers in water, the mean free path of bremsstrahlung and pair production processes, or the radiation length $X_{0}$, is about 36 cm [55]. The distance to the shower’s maximum is $X_{0}\ln(E_{0}/E_{c})$, where $E_{0}$ is the initial energy of the first particle starting the shower. The full length of the shower is several times larger [62, 63]. For hadronic showers, the mean free path for pion interactions is numerically somewhat larger than $X_{0}$, but the multiplicity of the particles in each generation is larger than in electromagnetic showers [63], leading to a comparable length for the showers. In Super-K, muon-induced electromagnetic and hadronic showers are less than 10 meters long even at the highest relevant energies, so they are often fully contained within the fiducial volume (when they are not, they can be treated as showers of lower energy).
Electromagnetic showers are quite narrow, with a characteristic (Molière) radius of $\sim$0.1 m in water [55], while hadronic showers, which have larger fluctuations at a given energy, have a characteristic radius of $\sim$1 m, about twice the hadronic interaction length (see the Appendices for details). While many of the secondary particles in both types of showers are collimated in the direction of the muon, lower-energy secondaries have a broader range of directions, which is important for shower reconstruction [32].
@@ II.4 Production of isotopes by showers
The total path length of the secondary particles in a shower is directly related to its initial energy because the number of secondary particles in a shower grows linearly with shower energy. The total energy in showers can be estimated from the total amount of Cherenkov light produced by the muon, minus the expectation from the non-showering energy loss components for the muon [67, 68, 32]. Therefore, the larger the light yield of a shower, the greater the cumulative path length of its secondary particles, which leads to a higher probability of producing an isotope.
@@ II.5 Super-K cuts on spallation isotopes
This improved likelihood cut was used in the latest solar neutrino analysis by Super-K [23]. In addition to that, they used a preselection cut that targets clusters of two or more isotopes produced by the same muon, which they referred to as a multiple spallation cut (Borexino used a similar multiple-isoptope approach to tag 11C isotopes [41]). Unlike the likelihood cut, the multiple spallation cut does not require muon reconstruction, but it does require the reconstruction of multiple isotope decays. Overall, the latest Super-K analysis achieved an efficiency of $\sim$90with a 10.5% deadtime [25], a significant improvement over earlier work by Super-K. For the data period with neutron tagging in pure water, they also applied cuts based on clouds of neutrons (two or more detected, which implies $\sim$10 times more produced) [50, 51, 23]. These cuts used geometric estimates of the shower size and position based on the multiplicity and locations of the neutron captures. The impact of these geometric cuts was limited by the low neutron statistics. We advise that it is better to estimate the shower size and position based on its abundant Cherenkov light. This neutron cloud technique further reduced the deadtime to 8.6% [23, 25].
As an aside, we advocate that the multiple spallation cut be replaced by a strong cut on muons with large total energy losses, as suggested in Ref. [31], which notes that muons with energy losses of 30 GeV or more produce $\simeq$60of isotopes and can be strongly cut while incurring only minimal deadtime. This approach has several advantages. First, it does not require multiple isotopes, so it applies to lower-energy showers. Second, it can be adapted to muons that are not vertically downgoing by using a cut at a maximum energy loss of $30~{}{\rm GeV}/32.2~{}{\rm m}=0.93~{}{\rm GeV/m}$. Third, as described next, it can be used effectively against large muon bundles. Full reconstruction of the muon(s) should not be necessary; it should be enough to localize the approximate entry and exit positions in the outer detector and to measure the total light produced.
Muon bundles are dangerous if $N$ muons are assumed to be $N-1$ or fewer, leading to a miscalculation of the distance of an isotope candidate from a muon track. Based on MACRO data [72, 73], double muons should occur at the $\sim$10level and higher-multiplicity muons at the $\sim$1level; the typical separation between muons is several meters. At present, Super-K attempts to reconstruct the positions and directions of simultaneous muons with even large multiplicities (up to 10 muons) [74]. For spallation work, this approach can be replaced with a simple cut based on the total energy. For example, four vertically downgoing muons would deposit a minimum total energy of $\sim$27 GeV, close to the value noted just above; cutting these events based on energy alone (or generalizing to energy loss per meter) would greatly simplify the work needed. Following reconstructions of not just the muon positions and directions, but also their energy-loss profiles [32], we expect that most of the cases with multiplicities of three muons could be distinguished from single showering muons and thus rejected. For the most relevant case of two muons, those with large separations can be treated independently, while those with small separations can be treated as a single, brighter muon.
@@ III.1 Average neutron yields
Neutrons, which are abundantly produced by muons in underground experiments, can cause a wide variety of difficult-to-reject detector backgrounds, not only through spallation decays. This is why neutrons are sometimes called “the rats of the lab.” There have been many measurements of neutron yields, but they have typically been reported as a function of the average energy of the muons at a given depth; see, e.g., Refs. [79, 49, 80, 44, 81]. These measurements have found a direct relationship between the average muon energy and the average neutron (or spallation) yields. While this is useful, the average rate of neutrons does not fully capture the underlying physics due to the large fluctuations in the energy losses of individual muons, and hence their neutron yields. Instead, it would be better to measure the neutron yield as a function of the parent shower energy, as this would put measurements at different depths on a common basis closely related to the underlying physics. As noted above, the shower energy can be determined from the measured total energy loss in the detector minus the predicted minimum energy loss. (At energies below the critical energy, it would be more accurate to say injected energy, as proper showers will not develop.) The neutron yield measurements would be even more informative if they were measured separately for hadronic and electromagnetic showers.
@@
In Appendix A, we give further details on the correlated yields of final-state neutrons, pions (and hence muons), and spallation isotopes, showing that they are all significantly higher for hadronic shower than electromagnetic showers (Fig. 13).
@@ IV.2 Neutron yields of hadronic showers are large
Figure 5 shows that hadronic showers almost always produce neutrons. As in Sec. III, we define $N_{n}$ as the number of low-energy neutrons left at the end of a shower; this is the same number that will undergo capture, for which the detection efficiency depends on the gadolinium concentration. The number of neutrons produced is often large (as shown above in Fig. 3). If we remake Fig. 5, instead requiring, for example, $N_{n}\geq 3$, the results only change slightly. The fraction of hadronic showers that do not produce any neutrons is only 3%. The logarithmic scale on the y-axis highlights the differences — which are statistically significant — in the energy spectra between the $N_{n}=0$ and $N_{n}\geq 1$ cases. The majority of the showers without neutrons are either low-energy hadronic showers or electromagnetic showers with small hadronic components, both of which are less efficient at producing isotopes anyway. These points mean that neutrons are extremely efficient at identifying hadronic showers, which is good.
[…]
@@ IV.4 Neutron-producing muons are rare
Figure 11 — another of our main results — shows the fraction of muons with neutron yields at or above a given value, assuming different tagging efficiencies. These values represent the fraction of the integral contributed by the bins in Fig. 7 above different $N_{n}$ thresholds. The sharp drop in the fraction of muons from $N_{n}=0$ to $N_{n}=1$ shows that about 89% of muons that cross Super-K produce no neutrons at all. This number results from the facts that each through-going muon produces on average one shower (as defined above), $\sim$97of these muon-induced showers are electromagnetic (Fig. 4), and $\sim$91of those electromagnetic showers do not produce neutrons (Fig. 6). Since the rate of all muons is $\sim$2 s-1, the rate of neutron-producing muons is only $\sim$0.2 s-1.
@@ IV.5 Recommendations for implementation
Figure 12: Deadtime and efficiency comparison for cylinder cuts based on shower energy (blue solid line) and neutron yields (dotted green line). The blue and green points represent different thresholds for $E_{\text{sh}}$ and $N_{n}$ with increments of 1 GeV and 1 neutron, respectively. The gray band highlights the efficiency improvement at $\sim$10deadtime.
@@
Figure 12 (before case) shows the deadtime and background-rejection efficiency for cylinder cuts where we vary the threshold for $E_{\text{sh}}$, meaning that showers above a certain energy are always rejected. These cuts can approximate (for a cut-off between 6–7 GeV) the performance of the current likelihood used by Super-K, which rejects $\sim$90of the backgrounds with $\sim$10deadtime.
@@ Acknowledgements
Data Availability
The data associated with the figures in this paper are available via Zenodo [85].
@@ Appendix A Final-state yields from showers
Figure 13: Numbers of neutron captures, muon decays, and isotopes left after hadronic and electromagnetic showers as a function of their energy. The differences between the hadronic and electromagnetic showers are large.
@@
Figure 13 shows the neutron, muon, and isotope yields in hadronic and electromagnetic showers as a function of the shower energy. The main feature is that the yields are approximately one order of magnitude lower in the electromagnetic case. In the hadronic case at the lowest energies, the results are significantly affected by the fact that half of the pions we inject are negative. In scintillator detectors, the muon and neutron yields are comparable, but the isotope yield is much higher, principally due to 11C (see Apeendix. C).
@@
Figure 15 shows the isotope yield distributions for muons at Super-K. This is related to Fig. 14 by a change of variables from $N_{n}$ to $N_{iso}$. The neutron yields of hadronic showers are $\sim$34 times larger than the isotope yields (see Fig. 13), so here the bin widths are changed to be $dN_{iso}=3$. The points we mentioned regarding the frequency of large neutron yields apply to isotope yields as well. For example, the daily muons with $\sim$1000 neutrons also produce $\sim$30 background isotope events, and the yearly muons with $\sim$8000 neutrons produce $\sim$240 isotopes. This large number of events can be utilized to test and calibrate the detector response to a supernova event.
Figure 14: Neutron yield distribution for muons at Super-K, similar to Fig. 7, but extending to larger neutron yields and using bin widths of of $dN_{n}=100$ instead of 1.
Figure 15: Isotope yield distribution for muons at Super-K, with bin widths of $dN_{iso}=3$.
@@ Appendix B Spatial profiles
Figure 16: Average longitudinal profiles of neutron captures, isotope production locations, and electron stopping points in 10-GeV hadronic and electromagnetic showers. The differences between hadronic and electromagnetic showers are small.
@@
Figure 16 shows the average longitudinal profiles of neutron captures, isotope production, and electron stopping points in 10-GeV electromagnetic and hadronic showers. Both have a longitudinal extent of $\sim$6 m, and the shapes of their longitudinal profiles have small differences. One of the main differences is that the neutron and isotope profiles in hadronic showers peak at a smaller distance. This is because the initial interactions of the injected pions can produce neutrons and isotopes.
Figure 17: Average lateral profiles of neutron captures, relevant isotope production locations, and electron stopping points in 10-GeV hadronic and electromagnetic showers. The differences between hadronic and electromagnetic showers are large.
Figure 17 shows the average lateral profiles of neutron captures, isotope production, and electron stopping points around the tracks of 10-GeV electromagnetic and hadronic showers. The area factor $2\pi rdr$ is included. Compared to the longitudinal profiles, the lateral profiles of electromagnetic showers and hadronic showers have large differences. In earlier work, we have shown that the lateral profiles of different isotopes are quite similar [30].
For electromagnetic showers, the lateral extent is characterized by the Molière radius ($\sim$10 cm in water). The majority of the shower energy is deposited within 0.5 m around the shower track, as seen in the bottom panel of Fig. 17. At 10 GeV, electromagnetic showers have an average neutron yield of $\sim$0.6. Neutrons at energies above tens of MeV can travel far away from the track because they undergo many scattering interactions before capturing, unlike thermal neutrons, but most of them still capture within $\sim$1.5 m from the shower. Isotope production in electromagnetic showers is mainly caused by these neutrons, so the isotopes are also contained within $\sim$1.5 m, but there is a larger peak at close distances to the shower because of isotopes produced by gamma rays directly or by neutrons that interact quickly.